OpenStudy (shes.radical):

Given the functions m(x) = 4x − 11 and n(x) = x − 10, solve m[n(x)] and select the correct answer below. (a) m[n(x)]=4x-51 (b) m[n(x)]=4x-29

OpenStudy (shes.radical):

(c) m[n(x)]=\[4x^2-51\] (d) m[n(x)]= \[4x^2-29\]

Nnesha (nnesha):

\[\large\textrm {m[n(x)] is same as m(x-10)} \]

Nnesha (nnesha):

substitute x for n(x) function which is x-10 into m(x) function

OpenStudy (shes.radical):

...what??

Nnesha (nnesha):

:P

Nnesha (nnesha):

n(x) =x-10 so they want u to replace x for x-10 into m(x) function so\[\huge\rm m(\color{Red}{x-10})=4\color{reD}{x}-11\] replace x with x-10

Nnesha (nnesha):

make sense ?

OpenStudy (shes.radical):

sorta..? ugh im so confused..now what?

OpenStudy (shes.radical):

so now its 4 (x-10) - 11 ?

OpenStudy (shes.radical):

its 4x-29! i get it now! lol, big help, thanks!

Nnesha (nnesha):

how??

Nnesha (nnesha):

distribute parentheses by 4 4(x-10))=? ?? distribute parentheses by 4

OpenStudy (shes.radical):

4(x-10)-11 = 4x- 40- 11 = 4x - 29

Nnesha (nnesha):

-40-11= ?

Nnesha (nnesha):

both are negative so you would add

OpenStudy (shes.radical):

oh..so then its 4x-51?

Nnesha (nnesha):

yes right

OpenStudy (shes.radical):

u rock..thank you!

Nnesha (nnesha):

my pleasure